Square root of 2x+5 - square root of x-2 =3
Answer by edjones(7569)

sqrt(2x+5)-sqrt(x-2)=3
2x+5-2sqrt(2x+5)(x-2)+x-2=9 Square each side
3x+3-2sqrt(2x+5)(x-2)=9
-2sqrt(2x+5)*sqrt(x-2)=-3x+6
4(2x+5)(x-2)=9x^2-36x+36 square each side.
8x^2+4x-40=9x^2-36x+36
x^2-40x+76=0 Subtract the left side from the right.
x=2, x=38
.
Ed